# Advanced safety filter based on SOS Control Barrier and Lyapunov Functions

### v2

@@ VI-B Including current reference tracking
For this subsection, the objective is to stabilize both the DC-link voltage to $1$, and the quadrature current to some (not yet a priori pre-specified) quadrature current reference $i_{q,ref}$. To render our design parametric in $i_{q,ref}$, we introduce a new state representing the quadrature current reference. Consequently, we augment the dynamics from 49 with $\dot{i}_{q,ref}=0$.
@@
The corresponding coordinates of the states and inputs are given by $x=\begin{bmatrix}\tilde{v}_{dc}&\tilde{i}_{d}&\tilde{i}_{q}&\tilde{i}_{q,ref}\end{bmatrix}^{T}$ and $u:=\begin{bmatrix}\tilde{m}_{d}&\tilde{m}_{q}\end{bmatrix}^{T}$, where $\tilde{i}_{q,ref}=i_{q,ref}$. The allowable set $\mathcal{X}_{a}$ is chosen through the following polynomials:
$w_{1}(x)$ $=((\tilde{v}_{dc}+0.3)/0.5)^{2}+(\tilde{i}_{d}/20)^{2}+(\tilde{i}_{q}/20)^{2}$
$+(\tilde{i}_{q,ref}/20)^{2}-1$
$w_{2}(x)$ $=(\tilde{v}_{dc}/20)^{2}+({\tilde{i}_{d}}/1.3)^{2}+({\tilde{i}_{q}}/1.3)^{2}$
$+(\tilde{i}_{q,ref}/20)^{2}-1.$
Correspondingly, a safe set $\mathcal{X}_{s}$ and the nominal region $\mathcal{X}_{n}$ can be computed for this 4-dimensional space. Figure 6(c) illustrates the safe set $\mathcal{X}_{s}$ and the nominal set $\mathcal{X}_{n}$ for $i_{q}=i_{q,ref}=0$. Table I summarizes the number of variables of the corresponding SDP for each part, as well as the number of constraints. The initialization procedure completed in 51 iterations, where each iteration took around 60 seconds. The main algorithm completed in 20 iterations, where each iteration took around 90 seconds. The degree of the polynomials are summarized in Table II.
The simulation shown in Figure 9 illustrates the reference tracking performance of our safety filter $u_{s}(x)$ compared to the basic safety filter $u_{\text{SF}}(x)$ introduced in the previous subsection. The safety filter, although slightly perturbing the legacy controller, guarantees finite-time convergence to the nominal set defined for a quadrature current reference $i_{q,ref}$ of $1$. However, the basic safety filter $u_{\text{SF}}(x)$ perturbs the legacy control action, leading to a much slower convergence. In general, as we approach the boundary of the safe set by increasing $i_{q,ref}$, the basic safety filter is more likely to intervene the legacy control action.
Figure 9: (Quadrature current reference tracking): The closed-loop simulation demonstrates the finite-time convergence guarantee of the advanced safety filter $u_{s}(x)$ towards the nominal region $\mathcal{X}_{n}$ for non-zero quadrature current reference $i_{q,ref}=1$p.u. Additionally, the advanced safety filter exhibits a faster convergence to the equilibrium compared to the basic safety filter $u_{\text{SF}}(x)$. The latter is more likely to disturb the legacy control action as the quadrature current reference $i_{q,ref}$ approaches the boundary of the safe set, resulting in a slower convergence.
@@ VII Conclusion
We successfully developed and implemented a novel safety filter approach that, to our knowledge, is the first of its kind to leave the existing legacy controller subject to input constraints unaffected. This innovation could lead to greater acceptance of safety filters in industrial environments and real-world scenarios, advancing our understanding of their implementation across a broader range of applications. Our next step is to integrate this safety filter with our industrial partners for the control of an advanced converter topology.
@@ VIII-A Proof of Lemma 1
Because the safe set $\mathcal{X}_{s}$ is compact and forward invariant, and the closed-loop dynamical system $\dot{x}=f(x)+G(x)u_{s}(x)$ is Lipschitz-continuous, there exists a unique trajectory $x(t)$ for all $x(0)\in\mathcal{X}_{s}$ and $t>0$. The CLF condition in 17 with $u:=u_{s}(x)$ can be written without dissipation rate as $\forall x\in\partial X_{n}$:
$\nabla V(x)^{T}\left(f(x)+G(x)u_{s}(x)\right)\leq 0.$ (58)
Using the tangent cone $T_{\mathcal{X}_{n}}(x)=\{z\mid\nabla V(x)^{T}z\leq 0\}$, the condition 58 can be formulated as $\forall x\in\partial X_{n}$:
@@
Hence, according to [18, Nagumo’s Theorem 4.7], $\mathcal{X}_{n}$ is forward invariant. This proves that for all $x(0)\in\mathcal{X}_{n}$, we can choose $T=0$. lt remains to show that for all other initial states, i.e., $x(0)\in\mathcal{X}_{t}$, there exists $T\geq 0$ such that $x(T)\in\mathcal{X}_{n}$, or equivalently $V\left(x(T)\right)\leq 0$.
As $\mathcal{X}_{t}\subseteq\mathcal{X}_{s}$ is compact (cf. Assumption 1) and the dissipative rate $d(x)$ defined in 18 is assumed to be strictly positive and Lipschitz-continuous on $\mathcal{X}_{t}$, the following minimum exists: $l=\underset{x\in\mathcal{X}_{t}}{\text{min}}\lambda(x)\left(V(x)+V_{0}\right)>0$. Therefore, the scalar function $V\left(x(t)\right)$ strictly decreases over time for all $x(0)\in\mathcal{X}_{t}$ due to the CLF condition in 17: $\dot{V}=\nabla V(x)\left(f(x)+G(x)u_{s}(x)\right)\leq-l<0$. As $\mathcal{X}_{s}$ is forward invariant, $V\left(x(t)\right)$ is upper bounded by $V\left(x(t)\right)\leq V\left(x(0)\right)-lt$ for all $x(0)\in\mathcal{X}_{t}$. Therefore, there exists some $T\leq V\left(x(0)\right)/l$ such that $V\left(x(T)\right)\leq 0$.
VIII-B Proof of Lemma 2
By dividing 26a and 26b by $s(x)>0$, we obtain
$\begin{split}\nabla B_{i}^{T}\left(f+Gu_{\text{sos}}\right)+\frac{\gamma_{0,i}}{s}B_{i}-\frac{\gamma_{1,i}^{T}}{s}B_{\mathcal{I}_{t}}&\leq 0\end{split}$ (59a)
$\begin{split}\nabla V^{T}\left(f+Gu_{\text{sos}}\right)+d+\frac{\gamma_{0,0}}{s}V-\frac{\gamma_{1,0}^{T}}{s}B_{\mathcal{I}_{t}}&\leq 0,\end{split}$ (59b)
@@ VIII-B Proof of Lemma 2
From 59a, by using the fact that $\gamma_{1,i}(x)\geq 0$, $s(x)>0$, and $B_{i}(x)=0$ and $B_{\mathcal{I}_{t}}(x)\leq 0$ on $\partial\mathcal{X}_{s}$, we infer, for each $i\in\mathcal{I}_{t}$, that for all $x\in\partial\mathcal{X}_{s}$ there exists $u:=u_{\text{sos}}(x)$ such that $\nabla B_{i}^{T}\left(f+Gu_{\text{sos}}\right)\leq 0$ whenever $B_{i}(x)=0$. Or, equivalently, for all $x\in\partial\mathcal{X}_{s}$ there exists $u:=u_{\text{sos}}(x)$ such that $\nabla B_{i}^{T}\left(f+Gu\right)\leq 0$ for all $i\in\text{Act}(x)$, as defined in 16. This proofs the CBF condition in 15 using the definition of the tangent cone under Assumption 2.
@@
From 59b, by using the fact that $\gamma_{0,0}(x)\geq 0$, $\gamma_{1,0}(x)\geq 0$, $s(x)>0$, and $V(x)\geq 0$ and $B_{\mathcal{I}_{t}}(x)\leq 0$ on $\mathcal{X}_{t}$, we deduce that for all $x\in\mathcal{X}_{t}$ there exists $u:=u_{\text{sos}}(x)$ such that $\nabla V^{T}\left(f+Gu\right)+\lambda\bigl{(}V-V_{0}\bigr{)}\leq 0$. This proofs the CLF condition 17.
@@
Finally, from 26c, by using the fact that $V(x)=0$ on $\partial\mathcal{X}_{n}$, we deduce that the inequality $\nabla V^{T}\left(f+Gu_{n}\right)\leq 0$ holds for all $x\in\partial\mathcal{X}_{n}$. This proofs the CBF condition in 19.
VIII-C Proof of Lemma 3
@@
We use the definition of a quadratic module in 4, to rewrite the SOS constraint in 28 as $\forall x\in\mathbb{R}^{n},\exists\gamma_{u}\in(\Sigma[x])^{t}$:
$-a_{i}u_{\text{SOS}}(x)+b_{i}+\frac{\gamma_{u}(x)^{T}}{s(x)}B_{\mathcal{I}_{t}}(x)\geq 0$
@@
for $i\in\mathcal{I}_{l}$, and $u_{\text{SOS}}(x)=p(x)/s(x)$. Since the $\gamma_{u}(x)\geq 0$, $s(x)>0$, and $B_{\mathcal{I}_{t}}(x)\leq 0$ on $\mathcal{X}_{s}$, we deduce, for all $i\in\mathcal{I}_{l}$, $-a_{i}u_{\text{SOS}}(x)+b_{i}\geq 0$ on $\mathcal{X}_{s}$. This concludes the first part of the proof.
@@
By substituting the variable $v$ in 30 with $u-u_{0}$, we can reformulate the SOS constraints by using the definition of a quadratic module as $\forall x\in\mathbb{R}^{n},\exists\gamma_{u,1}\in\Sigma[x,u],\gamma_{u,2}\in(\Sigma[x,u])^{t}$:
@@
$\begin{split}&-\left((u-u_{0})^{T}Q_{u}\left(u_{\text{SOS}}(x)-u_{0}\right)-u_{\text{max}}^{2}\right)\\ &\quad+\frac{\gamma_{u,1}(x,u)}{s(x)}f_{u}(x,u-u_{0})+\frac{\gamma_{u,2}(x,u)^{T}}{s(x)}B_{\mathcal{I}_{t}}\geq 0,\end{split}$ (60)
@@
where $x\in\mathbb{R}^{n}$ and $u\in\mathbb{R}^{m}$ are the variables of the polynomial expressions. Because of its positive semidefiniteness, $Q_{u}$ can be decomposed into $Q_{u}=B_{u}^{T}B_{u}$. Furthermore, since $\gamma_{u,1}(x,u)\geq 0$, $\gamma_{u,2}(x,u)\geq 0$, $s(x)>0$, $B_{\mathcal{I}_{t}}(x)\leq 0$ on $\mathcal{X}_{s}$, and $f_{u}(x,u-u_{0})\leq 0$ on $\mathcal{U}$, condition 60 can be reformulated as $\forall x\in\mathcal{X}_{s},\forall u\in\mathcal{U}$:
$(u-u_{0})^{T}B_{u}^{T}B_{u}(u_{\text{SOS}}(x)-u_{0})\leq u_{\text{max}}^{2}.$ (61)
For any $x\in\{x\in\mathcal{X}_{s}\mid B_{u}(u_{\text{SOS}}(x)-u_{0})=0\}$, we have $(u_{\text{SOS}}(x)-u_{0})^{T}B_{u}^{T}B_{u}(u_{\text{SOS}}(x)-u_{0})=0\leq u_{\text{max}}^{2}$. Hence, $u_{\text{SOS}}(x)\in\mathcal{U}$.
@@
$\begin{split}\left(B_{u}\lambda(u_{\text{SOS}}(x)-u_{0})\right)^{T}\left(B_{u}\lambda(u_{\text{SOS}}(x)-u_{0})\right)&=u_{\text{max}}^{2}.\end{split}$ (62)
Additionally, substituting $u$ by $\lambda(u_{\text{SOS}}(x)-u_{0})+u_{0}$ in condition 61 yields the condition that $\forall x\in\mathcal{X}_{s}$:
$\lambda(B_{u}(u_{\text{SOS}}(x)-u_{0}))^{T}(B_{u}(u_{\text{SOS}}(x)-u_{0}))\leq u_{\text{max}}^{2}.$ (63)
By comparing 62 and 63, we conclude that $\lambda$ must be greater than $1$. As a consequence, $(B_{u}(u_{\text{SOS}}(x)-u_{0}))^{T}(B_{u}(u_{\text{SOS}}(x)-u_{0}))\leq u_{\text{max}}^{2}$ holds for all $x\in\mathcal{X}_{s}$. This concludes that $u_{\text{SOS}}(x)\in\mathcal{U}$ for all $x\in\mathcal{X}_{s}$.
VIII-D Proof of Theorem 2

### v3

@@ VI-B Including current reference tracking
For this subsection, the objective is to stabilize both the DC-link voltage to 11, and the quadrature current to some (not yet a priori pre-specified) quadrature current reference $i_{q,ref}$. To render our design parametric in $i_{q,ref}$, we introduce a new state representing the quadrature current reference. Consequently, we augment the dynamics from 59 with $\dot{i}_{q,ref}=0$.
@@
The corresponding coordinates of the states and inputs are given by $x=\begin{bmatrix}\tilde{v}_{dc}&\tilde{i}_{d}&\tilde{i}_{q}&\tilde{i}_{q,ref}\end{bmatrix}^{\top}$ and $u:=\begin{bmatrix}\tilde{m}_{d}&\tilde{m}_{q}\end{bmatrix}^{\top}$, where $\tilde{i}_{q,ref}=i_{q,ref}$.
Correspondingly, a safe set $\mathcal{X}_{s}$ and the nominal region $\mathcal{X}_{n}$ can be computed for this 4-dimensional space. Figure 6(c) illustrates the safe set $\mathcal{X}_{s}$ and the nominal region $\mathcal{X}_{n}$ for $i_{q}=i_{q,ref}=0$. Table I summarizes the number of variables of the corresponding SDP for each part, as well as the number of constraints. The initialization procedure completed in 51 iterations, where each iteration took around 60 seconds. The main algorithm completed in 20 iterations, where each iteration took around 90 seconds.
@@ VII Conclusion
We successfully developed and implemented a novel safety filter approach that, to our knowledge, is the first of its kind to leave the existing legacy controller subject to input constraints unaffected. This innovation could lead to greater acceptance of safety filters in industrial environments and real-world scenarios, advancing their implementation across a broader range of applications. Unlike traditional safety filters, which often require manual tuning and lack formal guarantees on preserving the nominal controller, our approach provides an automated and principled alternative that ensures interpretability and minimal intervention by design. To support practical deployment, we also developed a software framework that automates the computation of the proposed safety filter. Future work will focus on a rigorous analysis of robustness to approximation errors in the legacy controller.
@@ VIII-A Proof of Lemma 1
Because the safe set $\mathcal{X}_{s}$ is compact and forward invariant, and the closed-loop dynamical system $\dot{x}=f(x)+G(x)u_{s}(x)$ is Lipschitz-continuous, there exists a unique trajectory $x(t)$ for all $x(0)\in\mathcal{X}_{s}$ and $t>0$. The CLF condition in 18 with $u:=u_{s}(x)$ can be written without dissipation rate as $\forall x\in\partial X_{n}$:
$\nabla V(x)^{\top}\left(f(x)+G(x)u_{s}(x)\right)\leq 0.$ (63)
Using the tangent cone $T_{\mathcal{X}_{n}}(x)=\{z\mid\nabla V(x)^{\top}z\leq 0\}$, the condition 63 can be formulated as $\forall x\in\partial X_{n}$:
@@
Hence, according to [5, Nagumo’s Theorem 4.7], $\mathcal{X}_{n}$ is forward invariant. This proves that for all $x(0)\in\mathcal{X}_{n}$, we can choose $T=0$. lt remains to show that for all other initial states, i.e., $x(0)\in\mathcal{X}_{t}$, there exists $T\geq 0$ such that $x(T)\in\mathcal{X}_{n}$, or equivalently $V\left(x(T)\right)\leq 0$.
As $\mathcal{X}_{t}\subseteq\mathcal{X}_{s}$ is compact (cf. Assumption 1) and the dissipative rate $d(x)$ defined in 19 is assumed to be strictly positive and Lipschitz-continuous on $\mathcal{X}_{t}$, the following minimum exists: $l=\underset{x\in\mathcal{X}_{t}}{\text{min}}\lambda(x)\left(V(x)+V_{0}\right)>0$. Therefore, the scalar function $V\left(x(t)\right)$ strictly decreases over time for all $x(0)\in\mathcal{X}_{t}$ due to the CLF condition in 18: $\dot{V}=\nabla V(x)\left(f(x)+G(x)u_{s}(x)\right)\leq-l<0$. As $\mathcal{X}_{s}$ is forward invariant, $V\left(x(t)\right)$ is upper bounded by $V\left(x(t)\right)\leq V\left(x(0)\right)-lt$ for all $x(0)\in\mathcal{X}_{t}$. Therefore, there exists some $T\leq V\left(x(0)\right)/l$ such that $V\left(x(T)\right)\leq 0$.
VIII-B Derivation of the SOS constraints 26
The objective of this section is to convert the empty-set condition 24b into the SOS constraint given in 26b. The derivation of 26a and 26c is analogous. Applying Putinar’s P-satz to 24b yields the SOS constraint
$-\nabla V^{T}f-d\in\textbf{M}(\nabla V^{T}G,-\nabla V^{T}G,V,-B_{\mathcal{I}_{t}}).$
By utilizing the definition of a quadratic module in 5, we can express this SOS constraint by the condition that there exist two SOS polynomials $p_{0,1},p_{0,2}\in(\Sigma[x])^{m}$ such that
$-\nabla V^{T}\left(f+G(p_{0,1}-p_{0,2})\right)-d\in\textbf{M}(V,-B_{\mathcal{I}_{t}}).$
By exploiting the fact that the set of SOS polynomials, when combined with the set of negative SOS polynomials, spans the entire set of polynomials (i.e., $\Sigma[x]-\Sigma[x]=R[x]$), we can reformulate the SOS constraint as the condition that there exists a polynomial $p_{0}\in(R[x])^{m}$ such that
$-\nabla V^{T}\left(f+Gp_{0}\right)-d\in\textbf{M}(V,-B_{\mathcal{I}_{t}}).$
Finally, we introduce an additional SOS polynomial $s_{0}(x)$ (cf. 8) and rewrite the SOS constraint as the condition that there exists $s_{0}\in\Sigma[x],p_{0}\in(R[x])^{m}$ such that
$-\nabla V^{T}\left(s_{0}f+Gp_{0}\right)-s_{0}d\in\textbf{M}(V,-B_{\mathcal{I}_{t}}).$
The introduction of the polynomial $s_{0}(x)$ increases the set of feasible solutions when truncating the degrees of the involved polynomials, as described in Section II-B.
VIII-C Proof of Lemma 2
By dividing 27a and 27b by $s(x)>0$, we obtain
$\begin{split}\nabla B_{i}^{\top}\left(f+Gu_{\text{sos}}\right)+\frac{\gamma_{0,i}}{s}B_{i}-\frac{\gamma_{1,i}^{\top}}{s}B_{\mathcal{I}_{n_{\!B}}}&\leq 0\end{split}$ (64a)
$\begin{split}\nabla V^{\top}\left(f+Gu_{\text{sos}}\right)+d+\frac{\gamma_{0,0}}{s}V-\frac{\gamma_{1,0}^{\top}}{s}B_{\mathcal{I}_{n_{\!B}}}&\leq 0,\end{split}$ (64b)
@@ VIII-C Proof of Lemma 2
From 64a, by using the fact that $\gamma_{1,i}(x)\geq 0$, $s(x)>0$, and $B_{i}(x)=0$ and $B_{\mathcal{I}_{n_{\!B}}}(x)\leq 0$ on $\partial\mathcal{X}_{s}$, we infer, for each $i\in\mathcal{I}_{n_{\!B}}$, that for all $x\in\partial\mathcal{X}_{s}$ there exists $u:=u_{\text{sos}}(x)$ such that $\nabla B_{i}^{\top}\left(f+Gu_{\text{sos}}\right)\leq 0$ whenever $B_{i}(x)=0$. Or, equivalently, for all $x\in\partial\mathcal{X}_{s}$ there exists $u:=u_{\text{sos}}(x)$ such that $\nabla B_{i}^{\top}\left(f+Gu\right)\leq 0$ for all $i\in\text{Act}(x)$, as defined in 17. This proofs the CBF condition in 16 using the definition of the tangent cone under Assumption 2.
@@
From 64b, by using the fact that $\gamma_{0,0}(x)\geq 0$, $\gamma_{1,0}(x)\geq 0$, $s(x)>0$, and $V(x)\geq 0$ and $B_{\mathcal{I}_{n_{\!B}}}(x)\leq 0$ on $\mathcal{X}_{t}$, we deduce that for all $x\in\mathcal{X}_{t}$ there exists $u:=u_{\text{sos}}(x)$ such that $\nabla V^{\top}\left(f+Gu\right)+\lambda\bigl(V-V_{0}\bigr)\leq 0$. This proofs the CLF condition 18.
@@
Finally, from 27c, by using the fact that $V(x)=0$ on $\partial\mathcal{X}_{n}$, we deduce that the inequality $\nabla V^{\top}\left(f+Gu_{n}\right)\leq 0$ holds for all $x\in\partial\mathcal{X}_{n}$. This proofs the CBF condition in 20.
VIII-D Proof of Lemma 3
@@
We use the definition of a quadratic module in 5, to rewrite the SOS constraint in 29 as $\forall x\in\mathbb{R}^{n},\exists\gamma_{u}\in(\Sigma[x])^{\top}$:
$-a_{i}u_{\text{SOS}}(x)+b_{i}+\frac{\gamma_{u}(x)^{\top}}{s(x)}B_{\mathcal{I}_{n_{\!B}}}(x)\geq 0$
@@
for $i\in\mathcal{I}_{l}$, and $u_{\text{SOS}}(x)=p(x)/s(x)$. Since the $\gamma_{u}(x)\geq 0$, $s(x)>0$, and $B_{\mathcal{I}_{n_{\!B}}}(x)\leq 0$ on $\mathcal{X}_{s}$, we deduce, for all $i\in\mathcal{I}_{l}$, $a_{i}u_{\text{SOS}}(x)+b_{i}\geq 0$ on $\mathcal{X}_{s}$. This concludes the first part of the proof.
@@
By substituting the variable $v$ in 31 with $u-u_{0}$, we can reformulate the SOS constraints by using the definition of a quadratic module as $\forall x\in\mathbb{R}^{n},\exists\gamma_{u,1}\in\Sigma[x,u],\gamma_{u,2}\in(\Sigma[x,u])^{\top}$:
@@
$\begin{split}&-\left((u-u_{0})^{\top}Q_{u}\left(u_{\text{SOS}}(x)-u_{0}\right)-u_{\text{max}}^{2}\right)\\ &\quad+\frac{\gamma_{u,1}(x,u)}{s(x)}f_{u}(x,u-u_{0})+\frac{\gamma_{u,2}(x,u)^{\top}}{s(x)}B_{\mathcal{I}_{n_{\!B}}}\geq 0,\end{split}$ (65)
@@
where $x\in\mathbb{R}^{n}$ and $u\in\mathbb{R}^{m}$ are the variables of the polynomial expressions. Because of its positive semidefiniteness, $Q_{u}$ can be decomposed into $Q_{u}=B_{u}^{\top}B_{u}$. Furthermore, since $\gamma_{u,1}(x,u)\geq 0$, $\gamma_{u,2}(x,u)\geq 0$, $s(x)>0$, $B_{\mathcal{I}_{n_{\!B}}}(x)\leq 0$ on $\mathcal{X}_{s}$, and $f_{u}(x,u-u_{0})\leq 0$ on $\mathcal{U}$, condition 65 can be reformulated as $\forall x\in\mathcal{X}_{s},\forall u\in\mathcal{U}$:
$(u-u_{0})^{\top}B_{u}^{\top}B_{u}(u_{\text{SOS}}(x)-u_{0})\leq u_{\text{max}}^{2}.$ (66)
For any $x\in\{x\in\mathcal{X}_{s}\mid B_{u}(u_{\text{SOS}}(x)-u_{0})=0\}$, we have $(u_{\text{SOS}}(x)-u_{0})^{\top}B_{u}^{\top}B_{u}(u_{\text{SOS}}(x)-u_{0})=0\leq u_{\text{max}}^{2}$. Hence, $u_{\text{SOS}}(x)\in\mathcal{U}$.
@@
$\begin{split}\left(B_{u}\lambda(u_{\text{SOS}}(x)-u_{0})\right)^{\top}\left(B_{u}\lambda(u_{\text{SOS}}(x)-u_{0})\right)&=u_{\text{max}}^{2}.\end{split}$ (67)
Additionally, substituting $u$ by $\lambda(u_{\text{SOS}}(x)-u_{0})+u_{0}$ in condition 66 yields the condition that $\forall x\in\mathcal{X}_{s}$:
$\lambda(B_{u}(u_{\text{SOS}}(x)-u_{0}))^{\top}(B_{u}(u_{\text{SOS}}(x)-u_{0}))\leq u_{\text{max}}^{2}.$ (68)
By comparing 67 and 68, we conclude that $\lambda$ must be greater than 11. As a consequence, $(B_{u}(u_{\text{SOS}}(x)-u_{0}))^{\top}(B_{u}(u_{\text{SOS}}(x)-u_{0}))\leq u_{\text{max}}^{2}$ holds for all $x\in\mathcal{X}_{s}$. This concludes that $u_{\text{SOS}}(x)\in\mathcal{U}$ for all $x\in\mathcal{X}_{s}$.
VIII-E Proof of Theorem 2